y = e^sin x tanh^-1 (4x^2)

find dy/dx


1
Expert's answer
2021-04-28T16:34:54-0400

For the function "y = e^{\sin x} \tanh^{-1} (4x^2)" let us find "\frac{dy}{dx}:"


"\frac{dy}{dx}=e^{\sin x}\cos x\tanh^{-1} (4x^2)+e^{\sin x}\frac{8x}{1-16x^4}."